A Average Senior high school annual cost of tuition fee for all private schools last year was php 43.700. a random sample of cost this year for 45 private schools indicated that the sample mean was 45,800 and a sample standard deviation was php5,600.at level of significance, is there sufficient evidence to conclude that the cost increased?

1
Expert's answer
2022-02-25T08:59:37-0500

The following null and alternative hypotheses need to be tested:

"H_0: \mu\le43700"

"H_1:\mu>43700"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05,df=n-1=44" degrees of freedom, and the critical value for a right-tailed test is "t_c = 1.68023."

The rejection region for this right-tailed test is "R = \{t: t > 1.68023\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{X}-\mu}{s/\sqrt{n}}=\dfrac{45800-43700}{5600/\sqrt{45}}=2.51558"

Since it is observed that "t = 2.51558 > t_c = 1.68023," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed "t=2.51558," "df=44" degrees of freedom is "p=0.007804," and since "p=0.007804<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than "43700," at the "\alpha = 0.05" significance level.